OpenStudy (howard-wolowitz):

Consider the number of questions incorrect from classmates on a quiz { 10, 11, 12, 13, 13, 13, 14, 15, 16, 16, 17, 18, 18, 19, 20 }. According to the Empirical Rule, 68% of the data should fall between what two numbers? A. 12 and 17 B. 11.98 and 18.02 C. 13.5 and 16.35 D. 10.20 and 12.92

OpenStudy (howard-wolowitz):

We need to find someone who is good at Empirical Rule, can you help me find someone?

OpenStudy (howard-wolowitz):

@igigighjkl @Study312 @JoyKat

OpenStudy (howard-wolowitz):

@Vickyisthesmartone123

OpenStudy (joykat):

what is it that u need he1p with?

OpenStudy (howard-wolowitz):

my question

OpenStudy (igigighjkl):

What grade is this.

OpenStudy (joykat):

empirical rule says 68 percent (roughly) of data has to lie within one standard deviation from the mean...

OpenStudy (joykat):

d0 u n0w what mean means?

OpenStudy (joykat):

mean = average of all incorrect answers (225/15 = 15)

OpenStudy (howard-wolowitz):

yes

OpenStudy (joykat):

standard deviation is the square r00t of the variance...variance is the expected value of the squared difference between the variable's realization and the variable's mean:

OpenStudy (joykat):

Variance = [(10-15)^2 + (11-15)^2 + (12-15)^2 + (13-15)^2 + (13-15)^2 + (13-15)^2 + (14-15)^2 + (15-15)^2 + (16-15)^2 + (16-15)^2 + (17-15)^2 + (18-15)^2 + (18-15)^2 + (19-15)^2 + (20-15)^2] / 15

OpenStudy (joykat):

Variance = [25 + 16 + 9 + 4 + 4 + 4 + 1 + 0 + 1 + 1 + 4 + 9 + 9 + 16 + 25] / 15 Variance = 128 / 15 = 8.533333333333333333

OpenStudy (joykat):

SD = square root of the variance = 2.921 (approx) therefore, since 68% of the data is within +/- 1 standard deviation of the mean, it is between (15 - 2.921) and (15 + 2.921) ....which is 12.079 and 17.921

OpenStudy (howard-wolowitz):

thank you for showing me this an teachi gme

OpenStudy (howard-wolowitz):

Excuse me! I'm doing a study guide, for which I have to learn and study

OpenStudy (vickyisthesmartone123):

@Howard-Wolowitz Sorry! Thought you were interested cause you seem stuck hehe Sorry I won't do it again! :(